can you answer WACC View Original Response O Unfiltered Response Show all attempts ABC Companys current share price is $22.9 and it is expectedABC Company also has preferred stock outstanding that pays a $1.85 per share fixed dividend. If this stock is currently priceABC Company has 5 million common shares outstanding and 2 million preferred shares outstanding, and its equity has a total boABC Company faces a 35% tax rate. Using the correct answers from the previous parts, what is the companys WACC? Enter your a
Answer #1

W ACC = ЕР *rd*(1 – t) + = *re + *rp

D = 30 million

E = 22.9 * 5 = 114.5 million

P = 27.2 * 2 = 54.4 million

rd = 6.8%

re = 11.44%

rp = 6.8%

30 W ACC = 2 114.5 * 0.068*(1 – 0.35) + +6320*0.1144+ 54.4 198.9 *0.068

W ACC = 0.006666666667 + 0.06585620915 +0.0185982906

WACC = 0.09112116642

WACC = 9.11%

Can you please upvote? Thank You :-)